maria receives 225 dollars annual interest income on two investments. if the amount she invested at 6 percent was 450 more than the amount she invested at 8 percent how much did she invest at 6 percent and at 8 percent
Answer by richwmiller(17219)

We don't know the total invested.
We know the difference in the two accounts.
We know that the interest for the two accounts is $225
0.06*x+0.08*y=225
We know that the account at 6% has $450 more.
x=450+y
We substitute for x
0.06*(450+y)+0.08*y=225
We multiply out
27+0.06y+0.08*y=225
We combine like terms.
0.14*y=198
Isolate y
y=$1414.28571 at 8%
x=450+y
Calculate x
x=$1864.28571 at 6%
Now,we know the total invested is: 3278.57143
Total invested $1864.28571+$1414.28571=$3278.57143
We check
0.06*1864.28571+0.08*1414.28571=225
111.857143+113.142857=225
225.0=225
Since this statement is TRUE and neither x nor y is negative all is well.
codeintmt